OpenStudy (-welp-):

How is sqrt20 not in simplified radical form?

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle \sqrt{20}=\sqrt{4\times 5}=\sqrt{4}\times \sqrt{5} }\)

OpenStudy (solomonzelman):

and from there you can simplify it yourself.

OpenStudy (solomonzelman):

(you know what a square root of a 4 is, don't you)

OpenStudy (-welp-):

Oh. Thanks!

OpenStudy (solomonzelman):

yw

OpenStudy (solomonzelman):

Also, other examples, \(\color{#000000 }{ \displaystyle \sqrt{72}=\sqrt{36\times 2}=\sqrt{36}\times \sqrt{2}=6\times \sqrt{2}=6\sqrt{2} }\) \(\tiny{\\[1.2em]}\) \(\color{#000000 }{ \displaystyle \sqrt{75}=\sqrt{25\times 3}=\sqrt{25}\times \sqrt{3}=5\times \sqrt{3}=5\sqrt{3} }\) \(\tiny{\\[1.2em]}\)